Spg 2019 (2192) -Quiz 4: Chapter 71 2018 ing 1:13:42 A laboratory in Arizona is interested in finding the mean chloride level for a healthy resident in the state. A random sample of 60 healthy residents has a mean chloride level of 102 mEq/L. If it is known that the chloride levels in healthy individuals residing in Arizona have a standard deviation of 45 mEq/L, find a 99% confidence interval for the true mean chioride level of all healthy Arizona residents. Then complete the table below. Carry your intermediate computations to at least three decimal places. Round your answers to one decimal place. (If necessary, consult a list of formulas) What is the lower limit of the 99% confidence interval? What is the upper limit of the 99% confidence interval?
Answer #1

lower = xbar - z *sd/sqrt(n)
= 102 - 2.576 * 45/sqrt(60)
= 87.0347
=87.0

upper = xbar + z * sd/sqrt(n)
= 102 + 2.576 * 45/sqrt(60)
=116.96520765
=117.0

Please rate